Tammie wants to buy a scooter that costs $3,500. Her mom will loan her the money at 5% annual interest as long as she makes monthly payments over the next 3 years. How much will Tammie's monthly payment be?

N

I/Y

PV

PMT

FV

Harrison wants to double his money in 10 years. What annual rate of return must he find in an investment to reach his goal?

N

10

I/Y

7.2%

PV

PMT

FV
Answer #1

1.

Loan Amount = $3,500

Time Period = 3 years

Interest Rate = 5%

Calculating Monthly Payment,

Using TVM Calculation,

PMT = [PV = -3,500, N = 36, FV = 0, I/Y = 0.05/12]

PMT = $104.90

Monthly Payment = $104.90

2.

Time Period = 10 years

PV = 1

FV = 2

PMT = 0

Calculating Interest Rate,

Using TVM Calculation,

I/Y = [PV = -1, FV = 2, PMT = 0, T = 10]

I/Y = 7.18%

Annual Interest Rate = 7.18%

Please "Like" if you find this useful and comment if you have any doubts.